# MIRAGE: Multimodal foundation model and benchmark for comprehensive retinal OCT image analysis

### v2

@@ 3 Discussion
While this work has comprehensively validated MIRAGE in several tasks, there are still limitations and challenges that need to be explored in future work. First, all the data used to develop MIRAGE is from a single center (Medical University of Vienna, Austria), and most of the samples are from patients with AMD, with smaller number of samples from other diseases, such as diabetes. Therefore, it is important to pursue a larger dataset by including retinal images across the world with a more diverse and realistic data distribution. Second, the pretraining of MIRAGE and the evaluation of its OCT diagnostic performance were limited to the central 2D B-scans of the OCT volumes. This was done, in line with previous work [33], to avoid the high computational cost of processing the entire 3D OCT derived from the use of ViT models. However, this is a limitation, as the use of 3D information is essential in clinical practice, and previous work [86, 87, 58, 88] has shown that it can significantly improve the performance of models in retinal image analysis tasks. Thus, research on developing efficient ways to incorporate 3D information into the pretraining and evaluation of FMs is an important direction for future work, that could ultimately lead to the development of more accurate FMs. Third, retinal layer segmentations for pretraining were generated using a graph-theoretic approach and image processing techniques[47, 48]. While this algorithm is OCT device-agnostic, and guarantees the topological correctness of the segmentations, its accurate performance is limited to healthy images. For this reason, layer segmentations are considered as weak labels or pseduo-labels. It is likely that the segmentation performance on pathological images could be improved by using more sophisticated segmentation algorithms, which could even incorporate the segmentation of pathological structures such as fluid pockets. Such segmentations would likely further improve the representation learning of MIRAGE and its downstream performance, especially on highly pathological segmentation datasets. In particular, we plan to investigate the use of topology-aware deep learning algorithms such as SD-LayerNet [89]. Fourth, while MIRAGE has shown strong performance in segmentation tasks, especially in the cross-dataset evaluation, it is slightly less performant than some task-specific state-of-the-art models [89, 90]. This is likely due to the fact that these methods are specifically designed for the segmentation task at hand, while MIRAGE was tuned using a general and straightforward semantic segmentation setup. This was done to minimize the impact of modules other than the pretrained encoder on the results, and to focus primarily on the evaluation of the pretraining strategy. Nevertheless, MIRAGE is not inherently incompatible with more sophisticated segmentation methods and architectures, and it is likely that its performance could be improved by using any of these innovations. Exploring this question represents an interesting area for future work. Fifth, while the linear probing and decoder-only fine-tuning strategies allowed us to effectively evaluate MIRAGE and SOTA FMs and achieve strong performance in most tasks, more sophisticated adapters or tuning strategies could improve downstream performance. A common example is Low-Rank Adapters (LoRA) [91], which enables the adaptation of large models to downstream tasks by tuning a small set of modules interleaved with the original architecture. This approach has been shown to improve the performance of large models in OCT medical image analysis [45]. In addition, it is very likely that fully fine-tuning the models could improve downstream classification performance if enough data is available [92, 33], as is the case for segmentation. Unfortunately, there are not as many large-scale datasets available for OCT classification as there are for other imaging modalities, such as natural images, and, despite successful advances in the field [], it is not guaranteed that full fine-tuning would not lead to overfitting on small datasets. Investigating how to adapt MIRAGE to maximize performance while maintaining generalizability represents an interesting area for future work. Finally, the current version of MIRAGE only uses image information from the OCT and SLO modalities, and pseudo-labels of retinal layers. The inclusion of color fundus photography (CFP), which is considered to be more informative and much more widely used than SLO (especially in developing countries), would make MIRAGE even more useful. In this study, it was not possible to include CFP since, to the best of our knowledge, no large dataset of paired OCT–CFP data is available (either publicly or in our clinic), and this pairing is essential for multimodal pretraining. This is because, unlike OCT and SLO images (which are commonly acquired together by modern OCT devices), CFP images are usually acquired using a separate imaging device or not at all when fundus is examined primarily with a slit lamp, making it difficult to collect paired OCT–CFP data. In addition, incorporating data such as clinical notes or demographic information during pretraining could provide further meaningful feedback to the model, ultimately improving representations. In this sense, we believe that MIRAGE could open the path towards more effective vision language models (VLMs) for OCT. While generalist large language models (LLMs) have demonstrated strong language capabilities for ophthalmology, they have also demonstrated limited vision capabilities for feature detection [93, 94]. Therefore, incorporating a robust and generalizable vision encoder such as MIRAGE could improve the feature detection and thus overall performance. These limitations and future directions represent interesting research opportunities in the field and areas for improvement of the proposed FM.
@@
In conclusion, we proposed MIRAGE, the first multimodal vision foundation model for OCT and SLO image analysis. To demonstrate its effectiveness and efficiency in adapting to various healthcare applications, we evaluated MIRAGE and other FMs on a newly proposed benchmark. Unlike previous evaluation approaches, our benchmark includes segmentation tasks in addition to diagnosis and staging, and is composed of a heterogeneous set of 19 tasks from 16 datasets. The results on our benchmark show significant performance improvements of MIRAGE over existing FMs in detecting and staging ocular diseases and segmenting retinal lesions and layers in OCT and SLO images. MIRAGE was found to be particularly robust to dataset shifts and generalizable to different types of tasks and modalities, with the largest differences over existing FMs found in OCT segmentation. In light of these results, we believe that MIRAGE can serve as an effective multimodal foundation model for retinal image analysis, with potential applications in research and clinical practice. In addition, we hope that our benchmark will help to better assess the capability of FMs for OCT, facilitating future comparisons and tracking progress in the field.
@@ 4.2 Benchmark datasets
The selected retinal disease classification datasets, listed in Table 9, contain images from both normal eyes and from patients exhibiting retinal signs of various diseases and conditions, including glaucoma, DME, early, intermediate, and advanced AMD ([e,i,a]AMD), RVO, geographic atrophy (GA), macular hole (MH), central serous retinopathy (CSR), epiretinal membrane (ERM), retinal artery occlusion (RAO), vitreomacular interface disease (VID), Stargardt disease, and CNV. The scans were acquired using a diverse set of OCT devices: Spectralis, Cirrus, Triton (Topcon, Tokyo, Japan), RTVue XR (Optovue, Fremont, USA) or Bioptigen (Leica Microsystems, Wetzlar, Germany). In addition, the datasets come from clinics in 6 different countries: USA, China, Austria, Iran, Russia, and India. As it was done for the pretraining dataset and in previous work [33], we restricted the analysis to the central B-scans of the OCT volumes.
@@
We also compiled five retinal layer and lesion segmentation datasets to benchmark our model on retinal lesion and layer segmentation tasks (see Table 10). In addition to the public datasets, we included a private dataset with SLO images for the segmentation of geographic atrophy (GA). These datasets cover a wide range of retinal lesions and layers, some of which are subsets or combinations of others. In particular, the following lesions are included: cystoid edema (cyst), pigment epithelial detachment (PED), subretinal fluid (SRF), intraretinal fluid (IRF), and GA. The datasets also include the following retinal layers: inner limiting membrane (ILM), retinal nerve fiber layer (RNFL), retinal pigment epithelium (RPE), retinal pigment epithelium-drusen complex (RPEDC), ganglion cell layer (GCL), inner plexiform layer (IPL), inner nuclear layer (INL), outer plexiform layer (OPL), outer nuclear layer (ONL), external limiting membrane (ELM), inner segment myoid (ISM), outer segment (OS), ganglion cell-inner plexiform layer complex (GCIPL), Bruch’s membrane (BM), choroid, and choroid-sclera interface (below BM). The scans were acquired with Cirrus, Spectralis, and Triton scanners at clinics in 5 different countries: USA, China, Austria, Croatia, and the Netherlands. In contrast to the classification datasets, all the B-scans in the segmentation datasets are used for training and evaluating the models. This was done following standard practices in the literature, where layer and lesion segmentation tasks are usually performed B-scan-wise, and not on the full volumes [95, 75, 76].
More details on all these datasets are listed in the Supplementary Note 2, with illustrative examples in Supplementary Figures 1 and 2. When training the models on the tasks listed above, we followed the standard train-test splits for the public datasets when available. In cases where no standard splits were provided, we performed a random split into training, validation, and test sets using patient and label stratification to ensure balanced representation.
@@ 4.3 MIRAGE approach
$\mathcal{L}=\mathcal{L}_{\scriptscriptstyle O}+\mathcal{L}_{\scriptscriptstyle N}+\mathcal{L}_{\scriptscriptstyle L}=\|\hat{x}_{\scriptscriptstyle O},x_{\scriptscriptstyle O}\|_{\scriptscriptstyle 2}+\|\hat{x}_{\scriptscriptstyle N},x_{\scriptscriptstyle N}\|_{\scriptscriptstyle 2}+\text{CE}(\hat{x}_{\scriptscriptstyle L},x_{\scriptscriptstyle L})\leavevmode\nobreak\ ,$ (1)
where the subscripts $O$, $N$, and $L$ denote OCT, SLO, and retinal layers pseudo-labels, and $\hat{x}$ and $x$ are the predicted and ground truth patches, respectively.
@@ 4.4 Network architecture
Figure 9: Overview of the ViT encoder. The linear projection layer projects the input patches into embeddings of dimension $d$, which are then passed through a stack of $L$ Transformer blocks containing $A$ attention heads.
@@ 4.5 Implementation details
All experiments were implemented using PyTorch [96] (pytorch.org) and the timm library (huggingface.co/docs/timm). For the evaluation, the scikit-learn (scikit-learn.org) library was used. Following previous work [33], we used the Vision Transformer (ViT) [29] architecture pretrained on ImageNet using MAE [18] as the backbone of our model. The main experiments were conducted using the ViT-Large architecture, while a version of MIRAGE based on the ViT-Base architecture was trained to study the impact of domain-specific multimodal pretraining and model capacity.
@@ 4.7 Segmentation evaluation procedure
To evaluate the accuracy of the predicted regions, we followed standard practices in the literature [104, 61, 63, 64, 42] and calculated the volume Dice coefficient, volume intersection over union (IoU), and 95th percentile of the Hausdorff distance (HD95). Dice coefficient is a measure of the overlap between the predicted and manual segmentations, with a value of 1 indicating perfect overlap and 0 indicating no overlap. It is calculated as the intersection of the predicted and manual segmentations divided by the sum of their volumes. IoU is a similar metric that is calculated as the intersection of the segmentations divided by their union. Finally, HD95 is a measure of the maximum distance between the predicted and manual segmentations, with the 95th percentile of the distances used to reduce the impact of outliers. HD95 is measured in pixels.
@@ Appendix 1 Supplementary Note 1: Extended results
In the following, we provide the extended tables of the experimental results, showing not only the average performance across all datasets, but also the results for each individual dataset. For the classification tasks, we report the area under the receiver operating characteristic curve (AUROC), average precision (AP), and balanced accuracy (BAcc). For the segmentation tasks, we report the Dice score, intersection over union (IoU), and 95th percentile of the Hausdorff distance (HD95) for all datasets except RETOUCH. For the RETOUCH dataset, we used the official evaluation server of the challenge to compute the metrics, which include the Dice score and absolute volume difference (AVD). To facilitate the comparison of the results, we highlight the best results in bold and underline the second best results. In addition, we have color-coded the best results according to the model. The color code is as follows (excluding models with no best results): MIRAGE (blue), RETFound [33] (red), DINOv2 [32] (green), SL-IN [29] (black), nnUNet [70] (magenta), MIRAGE–FFT (cyan), MultiMAE [23] (teal), MAE-OCT and MAE-SLO (salmon), MIRAGE-Base (blue-gray), OCT+Layers (lime), and TransUNet (dark gray). In all cases, the one-tailed Student’s $t$-test was used to assess the statistical significance of the performance difference between the best (in bold) and second best (underlined) models in each dataset. The Wilcoxon signed-rank test was used to assess the statistical significance of the performance differences across all datasets. Statistical significance is indicated by asterisks in the tables: *$p<0.05$, **$p<0.01$, ***$p<0.001$. In those cases where in-house datasets were used for the evaluation, we provide the average performance across both all the datasets and only public datasets (excluding in-house datasets).
@@ 2.1 Classification datasets
The Duke intermediate age-related macular degeneration (Duke iAMD) dataset includes 38,400 SD-OCT B-scans from 269 iAMD patients and 115 healthy subjects centered on a 5 mm diameter at the fovea. The images were acquired using a Bioptigen SD-OCT system (Research Triangle Park, NC) at four different clinics in the USA. Each volume has a retinal FOV of $\sim 6.7\times 6.7\leavevmode\nobreak\ \text{mm}^{2}$, and consists of 100 B-scans with dimensions of $512\times 1000$ pixels. In addition to the diagnosis, the dataset includes expert-annotated segmentation maps for three key retinal regions: the inner limiting membrane (ILM) to the inner retinal pigment epithelium (RPE) detachment complex (RPEDC), the inner RPEDC to the outer Bruch’s membrane (BM), and below the BM. For the classification evaluation, we focused on the detection of iAMD, while for the segmentation evaluation, we focused on the segmentation of all the aforementioned layers. Link: https://people.duke.edu/~sf59/RPEDC_Ophth_2013_dataset.htm
@@
This dataset originates from the Glaucoma grAding from Multi-Modality imAges (GAMMA) challenge [50], held in conjunction with the International Conference on Medical Image Computing and Computer-Assisted Intervention (MICCAI) 2023. It was provided by the Sun Yat-sen Ophthalmic Center, China, and contains pairs of color fundus and OCT images from randomly selected subjects with and without glaucoma. In particular, the public dataset includes 100 samples from the same number of patients. OCT volumes were acquired using a Topcon DRI OCT Triton device (Topcon, Tokyo, Japan). All volumes are centered on the macula with a FOV of $3\times 3$ mm. Each volume has 25 B-scans with dimensions of 992×512×992$\times$ 512$\times$992 × 512 × pixels. The dataset also includes glaucoma stages, foveal coordinates, and cup and optic disc segmentation masks. There are 26 samples with early glaucoma, 24 with moderate or advanced glaucoma, and 50 normal samples. In this study, we focus on the detection and staging of glaucoma. Link: https://gamma.grand-challenge.org/)
@@
The University of Minnesota (UMN) dataset, collected by the University of Minnesota Ophthalmology Clinic, USA, consists of 54 OCT volumes from 30 patients with DME and 24 patients with AMD. All scans were acquired using a Spectralis device centered on the macula. Each volume contains 25 B-scans with an image resolution of 496×1 0244961024496$\times$ 1\,024496 × 1 024 pixels. Link: https://people.ece.umn.edu/users/parhi/.DATA/

### v3

@@ 3 Discussion
While this work has comprehensively validated MIRAGE in several tasks, there are still limitations and challenges that need to be explored in future work. First, all the data used to develop MIRAGE is from a single center (Medical University of Vienna, Austria), and most of the samples are from patients with AMD, with smaller number of samples from other diseases, such as diabetes. Therefore, it is important to pursue a larger dataset by including retinal images across the world with a more diverse and realistic data distribution. Second, the pretraining of MIRAGE and the evaluation of its OCT diagnostic performance were limited to the central 2D B-scans of the OCT volumes. This was done, in line with previous work [33], to avoid the high computational cost of processing the entire 3D OCT derived from the use of ViT models. However, this is a limitation, as the use of 3D information is essential in clinical practice, and previous work [86, 87, 58, 88] has shown that it can significantly improve the performance of models in retinal image analysis tasks. Thus, research on developing efficient ways to incorporate 3D information into the pretraining and evaluation of FMs is an important direction for future work, that could ultimately lead to the development of more accurate FMs. Third, retinal layer segmentations for pretraining were generated using a graph-theoretic approach and image processing techniques[47, 48]. While this algorithm is OCT device-agnostic, and guarantees the topological correctness of the segmentations, its accurate performance is limited to healthy images. For this reason, layer segmentations are considered as weak labels or pseduo-labels. It is likely that the segmentation performance on pathological images could be improved by using more sophisticated segmentation algorithms, which could even incorporate the segmentation of pathological structures such as fluid pockets. Such segmentations would likely further improve the representation learning of MIRAGE and its downstream performance. In particular, we plan to investigate the use of topology-aware deep learning algorithms such as SD-LayerNet [89]. Fourth, while MIRAGE has shown strong performance in segmentation tasks, especially in the cross-dataset evaluation, it is slightly less performant than some task-specific SOTA models [89, 90]. This is likely due to the fact that these methods are specifically designed for the segmentation task at hand, while MIRAGE was tuned using a general and straightforward semantic segmentation setup. This was done to minimize the impact of modules other than the pretrained encoder on the results, and to focus primarily on the evaluation of the pretraining strategy. Nevertheless, MIRAGE is not inherently incompatible with more sophisticated segmentation methods and architectures, and it is likely that its performance could be improved by using any of these innovations. Exploring this question represents an interesting area for future work. Fifth, while the linear probing and decoder-only fine-tuning strategies allowed us to effectively evaluate MIRAGE and SOTA FMs and achieve strong performance in most tasks, more sophisticated adapters or tuning strategies could improve downstream performance. A common example is Low-Rank Adapters (LoRA) [91], which enables the adaptation of large models to downstream tasks by tuning a small set of modules interleaved with the original architecture. This approach has been shown to improve the performance of large models in OCT medical image analysis [45]. In addition, it is very likely that fully fine-tuning the models could improve downstream classification performance if enough data is available [92, 33], as is the case for segmentation. Unfortunately, there are not as many large-scale datasets available for OCT classification as there are for other imaging modalities, such as natural images, and, despite successful advances in the field [], it is not guaranteed that full fine-tuning would not lead to overfitting on small datasets. Investigating how to adapt MIRAGE to maximize performance while maintaining generalizability represents an interesting area for future work. Finally, the current version of MIRAGE only uses image information from the OCT and SLO modalities, and pseudo-labels of retinal layers. The inclusion of color fundus photography (CFP), which is considered to be more informative and much more widely used than SLO (especially in developing countries), would make MIRAGE even more useful. In this study, it was not possible to include CFP since, to the best of our knowledge, no large dataset of paired OCT–CFP data is available (either publicly or in our clinic), and this pairing is essential for multimodal pretraining. This is because, unlike OCT and SLO images (which are commonly acquired together by modern OCT devices), CFP images are either acquired using a separate imaging device or not acquired at all, when fundus is examined primarily with a slit lamp. This makes it difficult to collect paired OCT–CFP data. In addition, incorporating data such as clinical notes or demographic information during pretraining could provide further meaningful feedback to the model, ultimately improving representations. In this sense, we believe that MIRAGE could open the path towards more effective vision language models (VLMs) for OCT. Furthermore, while generalist large language models (LLMs) have demonstrated strong language capabilities for ophthalmology, they have also demonstrated limited vision capabilities for feature detection [93, 94]. Therefore, incorporating a robust and generalizable vision encoder such as MIRAGE could improve the feature detection and thus the overall performance. These limitations and future directions represent interesting research opportunities in the field and areas for improvement of the proposed FM.
@@
In conclusion, we proposed MIRAGE, the first multimodal vision foundation model for OCT and SLO image analysis. To demonstrate its effectiveness and efficiency in adapting to various healthcare applications, we evaluated MIRAGE and other FMs on a newly proposed benchmark. Unlike previous evaluation approaches, our benchmark includes segmentation tasks in addition to diagnosis and staging, and is composed of a heterogeneous set of 19 tasks from 16 datasets. The results on our benchmark show significant performance improvements of MIRAGE over existing FMs in detecting and staging ocular diseases and segmenting retinal lesions and layers in OCT and SLO images. In light of these results, we believe that MIRAGE can serve as an effective multimodal foundation model for retinal image analysis, with potential applications in research and clinical practice. In addition, we hope that our benchmark will help to better assess the capability of FMs for OCT, facilitating future comparisons and tracking progress in the field.
@@ 4.2 Benchmark datasets
The selected retinal disease classification datasets, listed in Table 9, contain images from both normal eyes and from patients exhibiting signs of various retinal diseases and conditions, including glaucoma, diabetic macular edema (DME), early, intermediate, and advanced AMD ([e,i,a]AMD), retinal vein occlusion (RVO), geographic atrophy (GA), macular hole (MH), central serous retinopathy (CSR), epiretinal membrane (ERM), retinal artery occlusion (RAO), vitreomacular interface disease (VID), Stargardt disease, and choroidal neovascularization (CNV). The scans were acquired using a diverse set of OCT devices: Spectralis, Cirrus, Triton (Topcon, Tokyo, Japan), RTVue XR (Optovue, Fremont, USA) or Bioptigen (Leica Microsystems, Wetzlar, Germany). In addition, the datasets come from clinics in 6 different countries: USA, China, Austria, Iran, Russia, and India. As was done for the pretraining dataset and in previous work [33], we restricted the analysis to the central B-scans of the OCT volumes.
@@
We also compiled five retinal layer and lesion segmentation datasets to benchmark our model on retinal lesion and layer segmentation tasks (see Table 10). In addition to the public datasets, we included a private dataset with SLO images for the segmentation of GA. These datasets cover a wide range of retinal lesions and layers, some of which are subsets or combinations of others. In particular, the following lesions are included: cystoid edema (cyst), pigment epithelial detachment (PED), subretinal fluid (SRF), intraretinal fluid (IRF), and GA. The datasets also include the following retinal layers: inner limiting membrane (ILM), retinal nerve fiber layer (RNFL), retinal pigment epithelium (RPE), retinal pigment epithelium-drusen complex (RPEDC), ganglion cell layer (GCL), inner plexiform layer (IPL), inner nuclear layer (INL), outer plexiform layer (OPL), outer nuclear layer (ONL), external limiting membrane (ELM), inner segment myoid (ISM), outer segment (OS), ganglion cell-inner plexiform layer complex (GCIPL), Bruch’s membrane (BM), choroid, and choroid-sclera interface (below BM). The scans were acquired with Cirrus, Spectralis, and Triton scanners at clinics in 5 different countries: USA, China, Austria, Croatia, and the Netherlands. In contrast to the classification datasets, all the B-scans in the segmentation datasets are used for training and evaluating the models. This was done following standard practices in the literature, where layer and lesion segmentation tasks are usually performed B-scan-wise, and not on the full volumes [95, 75, 76].
More details on all these datasets are listed in the Supplementary Note 2, with illustrative examples in Supplementary Figures 1 and 2. When training the models on the tasks listed above, we followed the standard train-test splits for the public datasets when it was available. In cases where no standard splits were provided, we performed a random split into training, validation, and test sets using patient and label stratification to ensure balanced representation in each set.
@@ 4.3 MIRAGE approach
$\mathcal{L}=\mathcal{L}_{\scriptscriptstyle O}+\mathcal{L}_{\scriptscriptstyle N}+\mathcal{L}_{\scriptscriptstyle L}=\|\hat{x}_{\scriptscriptstyle O},x_{\scriptscriptstyle O}\|_{\scriptscriptstyle 2}+\|\hat{x}_{\scriptscriptstyle N},x_{\scriptscriptstyle N}\|_{\scriptscriptstyle 2}+\text{CE}(\hat{x}_{\scriptscriptstyle L},x_{\scriptscriptstyle L})\penalty 10000\ ,$ (1)
where the subscripts $O$, NN, and $L$ denote OCT, SLO, and retinal layers pseudo-labels, and $\hat{x}$ and $x$ are the predicted and ground truth patches, respectively.
@@ 4.4 Network architecture
Figure 9: Overview of the ViT encoder. The linear projection layer projects the input patches into embeddings of dimension $d$, which are then passed through a stack of $L$ Transformer blocks containing AA attention heads.
@@ 4.5 Implementation details
All experiments were implemented using PyTorch [96] (pytorch.org) and the timm library (huggingface.co/docs/timm). For the evaluation, the scikit-learn (scikit-learn.org) library was used. Following previous work [33], we used the ViT [29] architecture pretrained on ImageNet using MAE [18] as the backbone of our model. The main experiments were conducted using the ViT-Large architecture, while a version of MIRAGE based on the ViT-Base architecture was trained to study the impact of domain-specific multimodal pretraining and model capacity.
@@ 4.7 Segmentation evaluation procedure
To evaluate the accuracy of the predicted segmentations, we followed standard practices in the literature [104, 61, 63, 64, 42] and calculated the volume Dice coefficient, volume intersection over union (IoU), and 95th percentile of the Hausdorff distance (HD95). Dice coefficient is a measure of the overlap between the predicted and manual segmentations, with a value of 1 indicating perfect overlap and 0 indicating no overlap. It is calculated as the intersection of the predicted and manual segmentations divided by the sum of their volumes. IoU is a similar metric, and it is calculated as the intersection of the segmentations divided by their union. Finally, HD is a measure of the maximum distance between the predicted and manual segmentations. The 95th percentile of the distances, HD95, is used in order to reduce the impact of outliers. This metric is measured in pixels.
@@ Appendix 1 Supplementary Note 1: Extended results
In the following, we provide the extended tables of the experimental results, showing not only the average performance across all datasets, but also the results for each individual dataset. For the classification tasks, we report the area under the receiver operating characteristic curve (AUROC), average precision (AP), and balanced accuracy (BAcc). For the segmentation tasks, we report the Dice score, intersection over union (IoU), and 95th percentile of the Hausdorff distance (HD95) for all datasets except RETOUCH. For the RETOUCH dataset, we used the official evaluation server of the challenge to compute the metrics, which include the Dice score and absolute volume difference (AVD). To facilitate the comparison of the results, we highlight the best results in bold and underline the second best results. In addition, we have color-coded the best results according to the model. The color code is as follows (excluding models with no best results): MIRAGE[-Large] (blue), RETFound [33] (red), DINOv2 [32] (green), SL-IN [29] (black), nnUNet [70] (magenta), MIRAGE–FFT (cyan), MultiMAE [23] (teal), MAE-OCT and MAE-SLO (salmon), MIRAGE-Base (blue-gray), OCT+Layers (lime), and TransUNet (dark gray). In all cases, the one-tailed Student’s $t$-test was used to assess the statistical significance of the performance difference between the best (in bold) and second best (underlined) models in each dataset. The Wilcoxon signed-rank test was used to assess the statistical significance of the performance differences across all datasets. Statistical significance is indicated by asterisks in the tables: *$p<0.05$, **$p<0.01$, ***$p<0.001$. In those cases where in-house datasets were used for the evaluation, we provide the average performance across both all the datasets and only public datasets (excluding in-house datasets).
@@ 2.1 Classification datasets
The Duke intermediate age-related macular degeneration (Duke iAMD) dataset includes 38,400 SD-OCT B-scans from 269 iAMD patients and 115 healthy subjects centered on a 5 mm diameter at the fovea. The images were acquired using a Bioptigen SD-OCT system (Research Triangle Park, NC) at four different clinics in the USA. Each volume has a retinal FOV of $\sim 6.7\times 6.7\penalty 10000\ \text{mm}^{2}$, and consists of 100 B-scans with dimensions of $512\times 1000$ pixels. In addition to the diagnosis, the dataset includes expert-annotated segmentation maps for three key retinal regions: the inner limiting membrane (ILM) to the inner retinal pigment epithelium (RPE) detachment complex (RPEDC), the inner RPEDC to the outer Bruch’s membrane (BM), and below the BM. For the classification evaluation, we focused on the detection of iAMD, while for the segmentation evaluation, we focused on the segmentation of all the aforementioned layers. Link: https://people.duke.edu/~sf59/RPEDC_Ophth_2013_dataset.htm
@@
This dataset originates from the Glaucoma grAding from Multi-Modality imAges (GAMMA) challenge [50], held in conjunction with the International Conference on Medical Image Computing and Computer-Assisted Intervention (MICCAI) 2023. It was provided by the Sun Yat-sen Ophthalmic Center, China, and contains pairs of color fundus and OCT images from randomly selected subjects with and without glaucoma. In particular, the public dataset includes 100 samples from the same number of patients. OCT volumes were acquired using a Topcon DRI OCT Triton device (Topcon, Tokyo, Japan). All volumes are centered on the macula with a FOV of $3\times 3$ mm. Each volume has 25 B-scans with dimensions of $992\times 512\times$ pixels. The dataset also includes glaucoma stages, foveal coordinates, and cup and optic disc segmentation masks. There are 26 samples with early glaucoma, 24 with moderate or advanced glaucoma, and 50 normal samples. In this study, we focus on the detection and staging of glaucoma. Link: https://gamma.grand-challenge.org/)
@@
The University of Minnesota (UMN) dataset, collected by the University of Minnesota Ophthalmology Clinic, USA, consists of 54 OCT volumes from 30 patients with DME and 24 patients with AMD. All scans were acquired using a Spectralis device centered on the macula. Each volume contains 25 B-scans with an image resolution of $496\times 1\,024$ pixels. Link: https://people.ece.umn.edu/users/parhi/.DATA/